# Online estimation of the inverse of the Hessian for stochastic optimization with application to universal stochastic Newton algorithms

### v1

@@ 6.1 Study on the largest eigenvalue of $A_{n}$
Proof of Proposition 6.2.
First of all, note that for all $n\geq 0$, $\lambda_{\min}\left(A_{n}\right)>0$. According to the definition of $A_{n}$, one has
$\lambda_{\min}(A_{n+1})$ $=\lambda_{\min}\left(A_{n}-\gamma_{n+1}\left(T_{n+1}A_{n}+A_{n}T_{n+1}^{T}-2I_{d}\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)$
$\geq\lambda_{\min}\left(I_{d}-\gamma_{n+1}T_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)A_{n}$
×$\quad\times\left(I_{d}-\gamma_{n+1}T_{n+1}^{T}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)$
$\quad+\left.\left(\gamma_{n+1}I_{d}-\gamma_{n+1}^{2}T_{n+1}A_{n}T_{n+1}^{T}\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)$
$\geq\lambda_{\min}\left(I_{d}-\gamma_{n+1}T_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)A_{n}$
×$\quad\times\left(I_{d}-\gamma_{n+1}T_{n+1}^{T}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)$
$\quad+\gamma_{n+1}-\gamma_{n+1}^{2}\beta_{n+1}^{2}\lambda_{\max}(A_{n})-\gamma_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}$
For all $h\in\mathbb{R}^{p}$, one has
Thus,
Therefore,
$\lambda_{\min}(A_{n+1})$ $\geq\left(1-\gamma_{n+1}\beta_{n+1}\right)^{2}\lambda_{\min}\left(A_{n}\right)+\gamma_{n+1}-\gamma_{n+1}^{2}\beta_{n+1}^{2}\lambda_{\max}(A_{n})$
$\quad-\gamma_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}.$
Note that $(\beta_{n}\gamma_{n})_{n}$ is a decreasing sequence and $\beta_{n}\gamma_{n}\leq 1/2$. Let $U_{n}$ and $V_{n}$ be sequences defined as
$U_{n}:=\sum_{k=1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{k}\beta_{k}\right)^{2}\gamma_{k+1}^{2}\beta_{k+1}^{2}\lambda_{\max}(A_{n}),$
and
$V_{n}:=\sum_{k=0+1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{k}\beta_{k}\right)^{2}\gamma_{k+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{k+1}}.$
One can prove by induction that
$\lambda_{\min}(A_{n})$ $\geq\lambda_{\min}\left(A_{0}\right)\prod_{k=1}^{n}\left(1-\gamma_{k}\beta_{k}\right)^{2}+\sum_{k=1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}\gamma_{k}-U_{n}-V_{n},$
with the convention $\prod_{j=n+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}=1$. One has
since $\beta<\gamma-1/2$. One also has
$\geq\frac{1-\left(1-\gamma_{1}\beta_{1}\right)^{2}}{2\beta_{n}}.$
Thus, $\lambda_{\min}(A_{n})\geq\frac{c_{1}}{\beta_{n}}+o\left(\frac{1}{\beta_{n}}\right)$ a.s with $c_{1}=1-\left(1-\gamma_{1}\beta_{1}\right)^{2}/2$, which implies that
$\frac{1}{\lambda_{\min}(A_{n})}=O\left(\beta_{n}\right)\quad a.s.$
We now prove the convergence of $\theta_{n}$. According to Taylor’s theorem, and using Assumption (A2), we obtain
@@
It suggests that $\lim\inf_{n}\left\|\nabla G(\theta_{n})\right\|=0$ almost surely and $\lim\inf_{n}K_{n}=0$ almost surely. As $K_{n}$ converges almost surely to a random variable, $G(\theta_{n})$ converges almost surely to $G(\theta)$. By local strict convexity,
@@
Rate of convergence of $\theta_{n}$. Recall that
@@
According to equation (6) in Godichon-Baggioni, (2021), there is a positive constant $c_{0}$ such that
$\nabla G(\theta_{n})^{T}A_{n,\tau}\nabla G(\theta_{n})\geq\lambda_{\min}(A_{n,\tau})\left\|\nabla G(\theta_{n})\right\|^{2}\geq c_{0}\lambda_{\min}(A_{n,\tau})\left(G(\theta_{n})-G(\theta)\right).$
@@
$+\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}(n+1)^{\eta}.$
Let $\widetilde{U}_{n}:=\left(\frac{n+1}{n}\right)^{\eta}\left(\left(1+\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}\right)-c_{0}\nu_{n+1}^{2}\lambda_{\min}(A_{n,\tau})\right)$, then
$\mathbb{E}\left[\widetilde{V}_{n+1}|\mathcal{F}_{n}\right]$ $\leq\widetilde{V}_{n}+\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}(n+1)^{\eta}+\widetilde{V}_{n}\mathbf{1}_{\widetilde{U}_{n}>1}.$
As $\nu+\beta<1$, one can easily verify that $\mathbf{1}_{\widetilde{U}_{n}>1}\xrightarrow{a.s.}0$. We can now apply the Robbins-Siegmund theorem. We therefore have $\left\|G(\theta_{n+1})-G(\theta)\right\|=o\left(n^{-\eta}\right)$ for all $\eta<2\gamma+2\nu-3$. Thanks to the local strong convexity of $G$, this leads to $\left\|\theta_{n}-\theta\right\|^{2}=o\left(n^{-\eta}\right)$ a.s. We are now able to apply Theorem 3.1 and one obtains
@@
Rate of convergence of $\theta_{n,\tau^{\prime}}$. For all non-negative integers $n$,
@@
Rate of convergence of $t^{\prime}_{n}\sum_{k=0}^{n}\log(k+1)^{\tau^{\prime}}H^{-1}\left(\nabla G(\theta_{k})-\nabla_{h}g(X_{k+1},\theta_{k})\right)$. Analogous to the proof of Theorem 3.1, one can check with the help of a law of large numbers for martingales that
@@
Rate of convergence of $t^{\prime}_{n}\sum_{k=0}^{n}\log(k+1)^{\tau^{\prime}}H^{-1}\left(\nabla G(\theta_{k})-H(\theta_{k}-\theta)\right)$. With the help of Assumption (A3) and since $\theta_{n}$ converges almost surely to $\theta$, $\left\|G(\theta_{n})-H(\theta_{n}-\theta)\right\|=O\left(\left\|\theta_{n}-\theta\right\|^{2}\right)$ a.s. Then, with the help of Thanks to equality (25) and since $\nu<1$, we can prove with the help of the equation (24) that
@@
Rate of convergence of $t^{\prime}_{n}\sum_{k=0}^{n}\ln(k+1)^{\tau^{\prime}}H^{-1}A_{k+1,\tau}^{-1}\frac{\theta_{k}-\theta-(\theta_{k+1}-\theta)}{\nu_{k+1}}$. One can easily check that
@@
With the help of an Abel’s transform and since $A_{n,\tau}$ converges almost surely to $H^{-1}$, one has
@@
Thus, as $\left\|\theta_{n}-\theta\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\nu}}\right)a.s.$ we have for all $\delta>0$
@@
6.5 Useful lemmas

### v3

@@ 6.4.1 Rate of convergence of $A_{n}$
The proof is given in Section 6.7. According to Taylor’s theorem, and using Assumption (A2), we obtain
@@
It suggests that $\lim\inf_{n}\left\|\nabla G(\theta_{n})\right\|=0$ almost surely and $\lim\inf_{n}K_{n}=0$ almost surely. As $K_{n}$ converges almost surely to a random variable, $G(\theta_{n})$ converges almost surely to $G(\theta)$. By local strong convexity,
@@
6.6.1 Rate of convergence of $\theta_{n}$.
Recall that
@@
Since the functional $G$ is locally strongly convex meaning that there is a neighborhood of $\theta$ where smallest eigenvalues of the Hessian are uniformly bounded below by a positive constant. That is, the Polyak-Lojasiewicz inequality is locally satisfied: there is $r_{0}>0$ and $c_{0}>0$ such that for all $h\in\mathcal{B}(\theta,r_{0})$,
$\|\nabla G(h)\|^{2}\geq c_{0}(G(h)-G(\theta)).$
Then, since $\mathbf{1}_{\|\theta_{t}-\theta\|\leq r_{0}}=1-\mathbf{1}_{\|\theta_{t}-\theta\|>r_{0}}$,
$\left\|\nabla F\left(\theta_{t}\right)\right\|^{2}$ $\geq c_{0}(G\left(\theta_{t}\right)-G(\theta))\mathbf{1}_{\|\theta_{t}-\theta\|\leq r_{0}}$
$=c_{0}(G\left(\theta_{t}\right)-G(\theta))-c_{0}(G\left(\theta_{t}\right)-G(\theta)))\mathbf{1}_{\|\theta_{t}-\theta\|>r_{0}}$
@@
$+\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}(n+1)^{\eta}+2^{\eta}c_{0}\nu_{n+1}\lambda_{\min}(A_{n,\tau})\mathbf{1}_{\|\theta_{n,\tau}-\theta\|>r_{0}}\widetilde{V}_{n}.$
Let $\widetilde{U}_{n}:=\left(\frac{n+1}{n}\right)^{\eta}\left(\left(1+\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}\right)-c_{0}\nu_{n+1}\lambda_{\min}(A_{n,\tau})\right)$, then
$\mathbb{E}\left[\widetilde{V}_{n+1}|\mathcal{F}_{n}\right]$ $\leq\widetilde{V}_{n}+\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}(n+1)^{\eta}+\widetilde{V}_{n}(\mathbf{1}_{\widetilde{U}_{n}>1}+2^{\eta}c_{0}\nu_{n+1}\lambda_{\min}(A_{n,\tau})\mathbf{1}_{\|\theta_{n,\tau}-\theta\|>r_{0}}).$
As $\nu+\beta<1$, one can easily verify that $\mathbf{1}_{\widetilde{U}_{n}>1}\xrightarrow{a.s.}0$ and $\mathbf{1}_{\|\theta_{n,\tau}-\theta\|>r_{0}}\xrightarrow{a.s.}0$. Then,
$\sum_{t\geq 1}\widetilde{V}_{n}(\mathbf{1}_{\widetilde{U}_{n}>1}+2^{\eta}c_{0}\nu_{n+1}\lambda_{\min}(A_{n,\tau})\mathbf{1}_{\|\theta_{n,\tau}-\theta\|>r_{0}})<+\infty\quad a.s.$
In addition, since $12\gamma+2\nu-\eta-2>$ 1, it comes
$\sum_{t\geq 1}\frac{L_{\nabla GC}}{2}\nu_{n+1}^{2}\lambda_{\max}(A_{n,\tau})^{2}(n+1)^{\eta}<+\infty\quad a.s.$
We can now apply the Robbins-Siegmund theorem. We therefore have $G(\theta_{n})-G(\theta)=o\left(n^{-\eta}\right)$ for all $\eta<2\gamma+2\nu-3$. Thanks to the local strong convexity of $G$, this leads to $\left\|\theta_{n}-\theta\right\|^{2}=o\left(n^{-\eta}\right)$ a.s. We are now able to apply Theorem 3.1 and one obtains
@@
6.6.2 Rate of convergence of $\theta_{n,\tau^{\prime}}$.
For all non-negative integer $n$,
@@
Rate of convergence of $t^{\prime}_{n}\sum_{k=0}^{n}\log(k+1)^{\tau^{\prime}}H^{-1}\left(\nabla G(\theta_{k})-\nabla_{h}g(X_{k+1},\theta_{k})\right)$. Analogous to the proof of Theorem 3.1, since $\left(\nabla G(\theta_{k})-\nabla_{h}g(X_{k+1},\theta_{k})\right)_{k\geq 0}$ is a sequence of martingale differences, one can check with the help of a law of large numbers for martingales that
@@
Rate of convergence of $t^{\prime}_{n}\sum_{k=0}^{n}\log(k+1)^{\tau^{\prime}}H^{-1}\left(\nabla G(\theta_{k})-H(\theta_{k}-\theta)\right)$. With the help of Assumption (A3) and since $\theta_{n}$ converges almost surely to $\theta$, $\left\|\nabla G(\theta_{n})-H(\theta_{n}-\theta)\right\|=O\left(\left\|\theta_{n}-\theta\right\|^{2}\right)$ a.s. Then, with the help of equality (29) and since $\nu<1$, we can prove with the help of equality (28) that
@@
Rate of convergence of $t^{\prime}_{n}\sum_{k=0}^{n}\ln(k+1)^{\tau^{\prime}}H^{-1}A_{k+1,\tau}^{-1}\frac{\theta_{k}-\theta-(\theta_{k+1}-\theta)}{\nu_{k+1}}$. First, remark that
@@
With the help of an Abel’s transform and since $A_{n,\tau}$ converges almost surely to $H^{-1}$, using equality (29) and performing calculations analogous to those used in deriving equality (27), one has
@@
Thus, as $\left\|\theta_{n}-\theta\right\|^{2}=o\left(\frac{\ln n^{1+\delta}}{n^{\nu}}\right)a.s.$ and since $A_{k+1}$ and $A_{k,\tau}$ converge to $H^{-1}$, we have for all $\delta>0$
@@
6.7 Proof of Proposition 6.2
Note that for all $n\geq 0$, $\lambda_{\min}\left(A_{n}\right)>0$. According to the definition of $A_{n}$, one has
$\lambda_{\min}(A_{n+1})$ $=\lambda_{\min}\left(A_{n}-\gamma_{n+1}\left(T_{n+1}A_{n}+A_{n}T_{n+1}^{T}-2I_{d}\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right){-\lambda_{\max}\left(\Pi_{n+1}^{\perp}\right)}$
$\geq\lambda_{\min}(\left(I_{d}-\gamma_{n+1}T_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)A_{n}\left(I_{d}-\gamma_{n+1}T_{n+1}^{T}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)$
$\quad+\left(2\gamma_{n+1}I_{d}-\gamma_{n+1}^{2}T_{n+1}A_{n}T_{n+1}^{T}\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}){-\lambda_{\max}\left(\Pi_{n+1}^{\perp}\right)}$
$\geq\lambda_{\min}\left(I_{d}-\gamma_{n+1}T_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right)A_{n}\left(I_{d}-\gamma_{n+1}T_{n+1}^{T}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}\right))$
$\quad+2\gamma_{n+1}-\gamma_{n+1}^{2}\beta_{n+1}^{2}\lambda_{\max}(A_{n})-2\gamma_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{n+1}}{-\lambda_{\max}\left(\Pi_{n+1}^{\perp}\right)}$
For all $h\in\mathbb{R}^{p}$, one has
Thus,
Therefore,
$\lambda_{\min}(A_{n+1})$ $\geq\left(1-\gamma_{n+1}\beta_{n+1}\right)^{2}\lambda_{\min}\left(A_{n}\right)+2\gamma_{n+1}-\gamma_{n+1}^{2}\beta_{n+1}^{2}\lambda_{\max}(A_{n})$
$\quad-2\gamma_{n+1}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}{-\lambda_{\max}\left(\Pi_{n+1}^{\perp}\right)}.$
Note that $(\beta_{n}\gamma_{n})_{n}$ is a decreasing sequence. Let $U_{n}$, $V_{n}$ and $W_{n}$ be sequences defined as
$U_{n}:=\sum_{k=1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}\gamma_{k}^{2}\beta_{k}^{2}\lambda_{\max}(A_{n}),$
$V_{n}:=2\sum_{k=1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}\gamma_{k}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\geq\beta_{k+1}}$
and
$W_{n}:=\sum_{k=1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}{\lambda_{\max}\left(\Pi_{n+1}^{\perp}\right)}$
One can prove by induction that
$\lambda_{\min}(A_{n})$ $\geq\lambda_{\min}\left(A_{0}\right)\prod_{k=1}^{n}\left(1-\gamma_{k}\beta_{k}\right)^{2}+2\sum_{k=1}^{n}\prod_{j=k+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}\gamma_{k}-U_{n}-V_{n}{-W_{n}},$
with the convention $\prod_{j=n+1}^{n}\left(1-\gamma_{j}\beta_{j}\right)^{2}=1$. One has since $\mathbf{1}_{\Pi_{n}^{\perp}\neq 0}$ converges almost surely to 00,
$W_{n}\leq\prod_{k=1}^{n}\left(1-\gamma_{k}\beta_{k}\right)^{2}\sum_{k=1}^{n}\prod_{j=1}^{k}\left(1-\gamma_{j}\beta_{j}\right)^{-2}\lambda_{\max}\left(\Pi_{k}^{\perp}\right)=O\left(\prod_{k=1}^{n}\left(1-\gamma_{k}\beta_{k}\right)^{2}\right)\quad a.s.$
and this term converges exponentially fast to 00. In addition
since $\beta<\gamma-1/2$. One also has
$\geq\frac{1-\left(1-\gamma_{1}\beta_{1}\right)^{2}}{2\beta_{n}}.$
Thus, $\lambda_{\min}(A_{n})\geq\frac{c_{1}}{\beta_{n}}+o\left(\frac{1}{\beta_{n}}\right)$ a.s with $c_{1}=1-\left(1-\gamma_{1}\beta_{1}\right)^{2}$, which implies that
$\frac{1}{\lambda_{\min}(A_{n})}=O\left(\beta_{n}\right)\quad a.s.$
Appendix A How to verify assumptions in practice?
A.1 Linear model
We consider the case of the linear model, i.e., where $(X,Y)$ lies in $\mathbb{R}^{d}\times\mathbb{R}$ and
$Y=X^{T}\theta+\epsilon,$
with $\epsilon$ independent of $X$ and $\theta\in\mathbb{R}^{d}$. The functional to minimize is defined for all $h\in\mathbb{R}^{d}$ by
$G(h)=\mathbb{E}\left[\frac{1}{2}\left(Y-X^{T}h\right)^{2}\right]=:\mathbb{E}\left[g(X,Y,h)\right].$
Then, for all $x,y\in\mathbb{R}^{d}\times\mathbb{R}$, the function $g$ is twice differentiable, with
$\nabla_{h}g(x,y,h)=-(y-x^{T}h)x,\quad\quad\text{and}\quad\quad\nabla_{h}^{2}g(x,h)=xx^{T}.$
If $X$ admits a moment of order $2$, then $G$ is twice differentiable and, for all $h\in\mathbb{R}^{d}$,
$\nabla^{2}G(h)=\mathbb{E}\left[XX^{T}\right],$
so that assumption (A3) is satisfied. Suppose also that $\mathbb{E}\left[XX^{T}\right]$ is positive, ensuring assumption (A2). Additionally, suppose that $X$ admits a moment of order $2q$, which satisfies Assumption (A4). Finally, suppose that $X$ (resp. $\epsilon$) admits a moment of order $8$ (resp. $4$), ensuring that assumption (A1) holds. Indeed,
$\mathbb{E}\left[\left\|\nabla_{h}g(X,Y,h)\right\|^{2}\right]$ $\leq\mathbb{E}\left[\|X\|^{2}\left|X^{T}(\theta-h)+\epsilon\right|^{2}\right]$
$\leq 2\mathbb{E}\left[\|X\|^{4}\right]\|\theta-h\|^{2}+2\mathbb{E}\left[\epsilon^{2}\right]\mathbb{E}\left[\|X\|^{2}\right].$
Then, since for all $h\in\mathbb{R}^{d}$, $\nabla^{2}G(h)=\mathbb{E}\left[XX^{T}\right]$ is positive, we have
$\|\theta-h\|^{2}\leq\frac{2}{\mu}\left(G(h)-G(\theta)\right),$
where $\mu=\lambda_{\min}\left(\mathbb{E}\left[XX^{T}\right]\right)$, so that
$\mathbb{E}\left[\nabla_{h}g(X,Y,h)^{2}\right]\leq 2\mathbb{E}\left[\|X\|^{2}\right]\frac{2}{\mu}(G(h)-G(\theta))+2\mathbb{E}\left[\epsilon^{2}\right]\mathbb{E}\left[\|X\|^{2}\right].$
In addition,
$\mathbb{E}\left[\nabla_{h}g(X,Y,h)^{4}\right]$ $\leq\mathbb{E}\left[\|X\|^{4}\left|X^{T}(\theta-h)+\epsilon\right|^{4}\right]$
$\leq 8\mathbb{E}\left[\|X\|^{8}\right]\|\theta-h\|^{4}+8\mathbb{E}\left[\epsilon^{4}\right]\mathbb{E}\left[\|X\|^{4}\right],$
and Assumption (A1b) is satisfied taking any $R>0$.
Summary:
If $\epsilon$ admits a moment of order $4$, if $X$ admits a moment of order $\max\{8,2q\}$, and if $\mathbb{E}\left[XX^{T}\right]$ is positive, then the assumptions are satisfied.
A.2 Logistic regression
Let us consider the case of logistic regression, where $(X,Y)\in\mathbb{R}^{d}\times\{0,1\}$ and the conditional distribution of $Y$ given $X$ follows a Bernoulli law:
$Y|X\sim\mathcal{B}(\pi(\theta^{T}X)),$
where $\pi$ is the sigmoid function, defined for all $z\in\mathbb{R}$ as
$\pi(z)=\frac{e^{z}}{1+e^{z}}.$
The objective functional to minimize is given, for all $h\in\mathbb{R}^{d}$, by
$G(h)=\mathbb{E}\left[\log\left(1+\exp\left(h^{T}X\right)\right)-YX^{T}h\right]=:\mathbb{E}\left[g(X,Y,h)\right].$
For all $(x,y)\in\mathbb{R}^{d}\times\{0,1\}$, the function $g(x,y,\cdot)$ is twice differentiable, with
$\nabla_{h}g(x,y,h)=(\pi(x^{T}h)-y)x$
and
$\nabla_{h}^{2}g(x,y,h)=\pi(x^{T}h)(1-\pi(x^{T}h))xx^{T}.$
If $X$ has a finite fourth-order moment, Assumption (A1) holds, i.e.,
$\mathbb{E}\left[\left\|\nabla_{h}g(X,Y,h)\right\|^{4}\right]\leq\mathbb{E}\left[\left\|X\right\|^{4}\right].$
Furthermore, $G$ is also twice differentiable, with
$\nabla^{2}G(h)=\mathbb{E}\left[\pi(X^{T}h)(1-\pi(X^{T}h))XX^{T}\right],$
and
$\left\|\nabla^{2}G(h)\right\|_{op}\leq\frac{1}{4}\mathbb{E}\left[\|X\|^{2}\right].$
If $\nabla^{2}G(\theta)$ is positive, a condition satisfied, for instance, when $X$ follows an elliptic distribution (see Gadat and Panloup, (2017)), then Assumption (A2) is verified. Additionally, if $X$ has a finite third-order moment, Assumption (A3) holds (see Lemma 6.2 in Bercu et al., 2020b ):
$\left\|\nabla^{2}G(h)-\nabla^{2}G(\theta)\right\|\leq\frac{1}{12\sqrt{3}}\mathbb{E}\left[\|X\|^{3}\right]\|h-\theta\|.$
Finally, Assumption (A4) is satisfied if $X$ has a finite $2q$-th moment, since
$\mathbb{E}\left[\left\|\nabla_{h}^{2}g(X,Y,h)\right\|_{op}^{q}\right]\leq\frac{1}{4^{q}}\mathbb{E}\left[\|X\|^{2q}\right].$
Summary:
If $X$ admits a moment of order $\max\{3,2q\}$ and if $\nabla^{2}G(\theta)$ is positive (which is satisified if $X$ follows an elliptics distribution for instance), then the assumptions are satisfied.
A.3 Estimation of the median
Let us consider a continuous random vector $X$ in $\mathbb{R}^{d}$ satisfying the following two standard assumptions (see Cardot et al., (2013) among others):
  • The distribution of $X$ is not concentrated on a single straight line: for all $h\in\mathbb{R}^{d}$, there exists $h^{\prime}\in\mathbb{R}^{d}$ such that $h^{T}h^{\prime}=0$ and
    $\mathbb{V}\left[X^{T}h^{\prime}\right]>0.$
  • The distribution of $X$ is not concentrated around single points: there exists a constant $C$ such that for all $h\in\mathbb{R}^{d}$,
    $\mathbb{E}\left[\frac{1}{\|X-h\|^{2q}}\right]\leq C.$
This latter assumption is closely related to small-ball probabilities and is not restrictive as long as $d\geq 2q+2$ (see Cardot et al., (2013) among others). Under these conditions, the geometric median of $X$ is the unique minimizer of the functional defined for all $h\in\mathbb{R}^{d}$ by
$G(h):=\mathbb{E}\left[\|X-h\|-\|X\|\right]=:\mathbb{E}\left[g(X,h)\right].$
For all $h\in\mathbb{R}^{d}$ and $x\neq h$, the function $g(x,\cdot)$ is twice differentiable at $h$, with
$\nabla_{h}g(x,h)=-\frac{x-h}{\|x-h\|},\quad\quad\text{and}\quad\quad\nabla_{h}^{2}g(x,h)=\frac{1}{\|x-h\|}\left(I_{d}-\frac{(x-h)(x-h)^{T}}{\|x-h\|^{2}}\right).$
Thus, Assumption (A1) holds with
$\mathbb{E}\left[\left\|\nabla_{h}g(X,h)\right\|^{4}\right]=1.$
Furthermore, Assumption (A2) is satisfied since
$\nabla^{2}G(h)=\mathbb{E}\left[\frac{1}{\|X-h\|}\left(I_{d}-\frac{(X-h)(X-h)^{T}}{\|X-h\|^{2}}\right)\right],$
and
$\left\|\nabla^{2}G(h)\right\|_{op}\leq\mathbb{E}\left[\frac{1}{\|X-h\|}\right]\leq C^{1/2q}.$
The positivity of $\nabla^{2}G(m)$ (where $m$ is the median of $X$) is established by Proposition 2.1 in Cardot et al., (2013). Moreover, Assumption (A4) is also satisfied, and Assumption (A3) is verified in Lemma 3.3 of Cardot et al., (2013).
Summary:
If the distribution of the continuous random vector $X$ is not concentrated along a single straight line nor around single points, then all assumptions are satisfied.
A.4 Estimation of $p$-means
Let us now consider the estimation of the $p$-mean of a continuous random vector $X$ in $\mathbb{R}^{d}$, where $p\in(1,2)$. The $p$-mean $m$ of $X$ is defined as the minimizer of the functional defined for all $h\in\mathbb{R}^{d}$
$G(h)=\mathbb{E}\left[\left\|X-h\right\|^{p}\right]=:\mathbb{E}\left[g(X,h)\right].$
For any $x$, the function $g(x,.)$ is differentiable, and for any $h$, and for all $x\neq h$, it is twice differentiable at $h$, with
$\nabla_{h}g(x,h)$ $=-(x-h)\|x-h\|^{p-2},$
and
$\nabla_{h}^{2}g(x,h)$ $=\|x-h\|^{p-2}\left(I_{d}-(2-p)\frac{(x-h)(x-h)^{T}}{\|x-h\|^{2}}\right).$
Since $p\in(1,2)$, we have $2(p-1)<p$. Denoting by $m_{p}$ the $p$-mean of $X$, we obtain
$\mathbb{E}\left[\left\|\nabla_{h}g(X,h)\right\|^{2}\right]=\mathbb{E}\left[\left\|X-h\right\|^{2(p-1)}\right]\ \leq 1+\mathbb{E}\left[\left\|X-h\right\|^{p}\right]\ \leq 1+G(m_{p})+G(h)-G(m_{p}).$
Moreover,
$\mathbb{E}\left[\left\|\nabla_{h}g(X,h)\right\|^{4}\right]$ $=\mathbb{E}\left[\left\|X-h\right\|^{4(p-1)}\right]$
$\leq 2^{4p-5}\mathbb{E}\left[\left\|X-m_{p}\right\|^{4(p-1)}\right]+2^{4p-5}\left\|m_{p}-h\right\|^{4(p-1)}.$
Thus, Assumption (A1) holds for any $R>0$. Now, suppose that the distribution of $X$ is not concentrated around single points, i.e., there exists a constant $C_{q}$ such that for all $h\in\mathbb{R}^{d}$,
$\mathbb{E}\left[\frac{1}{\|X-h\|^{q(2-p)}}\right]\leq C_{q}.$
As in the case of the median, this assumption is closely related to small-ball probabilities and is non-restrictive as long as $d\geq 2q+2$. Under this assumption, Assumption (A4) is satisfied with
$\mathbb{E}\left[\left\|\nabla_{h}^{2}g(X,h)\right\|_{op}^{q}\right]\leq(2-p)^{q}\mathbb{E}\left[\frac{1}{\|X-h\|^{q(2-p)}}\right]\leq(2-p)^{q}C_{q}.$
Furthermore, analogous to the median, one can verify that Assumption (A3) holds (for instance, by adapting the proof of Lemma 5.1 in Cardot et al., (2017)). Finally, we have
$\nabla^{2}G(h)=\mathbb{E}\left[\|X-h\|^{p-2}\left(I_{d}-(2-p)\frac{(X-h)(X-h)^{T}}{\|X-h\|^{2}}\right)\right]$
and
$\lambda_{\min}\left(\nabla^{2}G\left(m_{p}\right)\right)\geq(p-1)\mathbb{E}\left[\frac{1}{\left\|X-m_{p}\right\|^{2-p}}\right]>0,$
thus satisfying Assumption (A2).
Summary:
If the distribution of $X$ is not concentrated around single points and if $X$ has a finite moment of order $4(p-1)$, then all the required assumptions are satisfied.
Appendix B Useful lemmas